OpenStudy (anonymous):

he motion of a set of particles moving along the x-axis is governed by the differential equation; dt/dx t^3-x^3,where x(t) denotes the position at time t of the particle. Show that the acceleration of a particle is given by d^2x/dt^2=3t^2-3t^33x^2 +3x^5 .

OpenStudy (experimentx):

'dt/dx t^3-x^3' is not an equation ... it's an expression, can you revise the question??

OpenStudy (anonymous):

its dt/dx= t^3-x^3 sorry!!

OpenStudy (experimentx):

dx/dt = 1/(t^3-x^3) \[ \frac{d^2x}{dt^2} = \frac{-1}{(t^3 - x^3)^2} (3t^2 - 3x^2 \frac{dx}{dt}) \\ =\frac{-1}{(t^3 - x^3)^2} (3t^2 - 3x^2 (t^3 - x^3))\] simplify ... simplify

OpenStudy (experimentx):

I think you meant dx/dt= t^3-x^3 rather than dt/dx= t^3-x^3

OpenStudy (anonymous):

yes I did mean dx/dt! but y is dx/dt= 1/(t^3-X^3)^2

OpenStudy (experimentx):

Oh ... just cut that -1/(t^3-x^3)^2 part ... you will have your answer!!

OpenStudy (anonymous):

i just dont understand how you get 3t^2-3x^2?

OpenStudy (anonymous):

(3t^2−3x^2dxdt) this part

OpenStudy (experimentx):

yes ...and put dx/dt = (t^3 - x^3)

OpenStudy (anonymous):

ok, but where is this part coming from (3t^2−3x^2dxdt)

OpenStudy (anonymous):

sorry just a little confused

OpenStudy (experimentx):

differentiation ... dt^3/dt - dx^3/dt

OpenStudy (anonymous):

wouldnt that give you 3t^2

OpenStudy (experimentx):

yes ... and the other would give your 3x^2 dx/dt

OpenStudy (anonymous):

sorry i have been away from calculus for a while, how does dx^3/dt = 3x^2? wouldn't it be equal to 0?

OpenStudy (experimentx):

if x is not a function of t, then it would be 0, but since x is a function of t, this is not zero!!

OpenStudy (experimentx):

i mean x is x(t) i.e. x depends on t ... if x were independent of t then it would have been zero

OpenStudy (anonymous):

oh ok i understand! Thank you!!

OpenStudy (experimentx):

yw